# Bayesian Inference in Epidemic Modelling: A Beginner’s Guide

### 16 Mar 2026 v1

@@ A Beginner’s Guide
When we build an epidemiological model, say, for COVID-19, influenza or Ebola, we face a fundamental challenge: we do not know the model’s parameters. How infectious is the pathogen? How quickly do people recover? These are the numbers that drive the model, yet they must be inferred from noisy, unobserved or incomplete data.
@@
Asks: “What is $\hat{\beta}$?” Asks: “What is $P(\beta\mid\text{data})$?”
@@
The SIR model is the workhorse of compartmental epidemiology. It divides a population of size $N$ into three groups:
@@
  • Susceptible $S(t)$: individuals who have never been infected and have no immunity. They are at risk of contracting the disease upon contact with an infectious person. At the start of an outbreak, nearly the entire population sits in this compartment.
  • Infectious $I(t)$: individuals who are currently infected and capable of transmitting the pathogen to susceptible individuals. The size of this compartment at any moment determines the force of infection driving the epidemic. In the basic SIR model, we assume that all infectious individuals are equally and immediately infectious, no latency period.
  • Recovered $R(t)$: individuals who have cleared the infection and are assumed to have acquired permanent immunity. They can neither be reinfected nor transmit the disease. In some diseases this is a reasonable approximation (e.g. measles); in others (e.g. influenza, COVID-19) immunity wanes over time, requiring extended models such as SIRS or SEIRS.
@@ 2.1  Solving the ODE: From Equations to Epidemic Curves
The idea is identical to Euler, follow the slope, step by step, but with a much more accurate estimate of which slope to follow. The scipy function odeint uses an adaptive version of this internally, automatically shrinking $\Delta t$ in regions where the solution changes rapidly (e.g. near the epidemic peak) and enlarging it where the solution is flat.

### 17 Mar 2026 v2

@@ A Beginner’s Guide
When we build an epidemiological model, say, for COVID-19, influenza or Ebola, we face a fundamental challenge: we do not know the model’s parameters $\theta$. How infectious is the pathogen? How quickly do people recover? These are the numbers that drive the model, yet they must be inferred from noisy, unobserved or incomplete data.
@@
Asks: “What is $\hat{\theta}$?” Asks: “What is $P(\theta\mid\text{data})$?”
@@
The Susceptible-Infected-Recovered (SIR) model is the workhorse of compartmental epidemiology. It divides a population of size $N$ into three groups:
@@
  • Susceptible $S(t)$: individuals who have never been infected at time $t$ and have no immunity. They are at risk of contracting the disease upon contact with an infectious person. At the start of an outbreak, nearly the entire population sits in this compartment.
  • Infectious $I(t)$: individuals who are currently infected at time $t$ and capable of transmitting the pathogen to susceptible individuals. The size of this compartment at any moment determines the force of infection driving the epidemic. In the basic SIR model, we assume that all infectious individuals are equally and immediately infectious, no latency period.
  • Recovered $R(t)$: individuals who have cleared the infection at time $t$ and are assumed to have acquired permanent immunity. They can neither be reinfected nor transmit the disease. In some diseases this is a reasonable approximation (e.g. measles); in others (e.g. influenza, COVID-19) immunity wanes over time, requiring extended models such as SIRS or SEIRS.
@@ 2.1  Solving the ODE: From Equations to Epidemic Curves
The idea is identical to Euler, follow the slope, step by step, but with a much more accurate estimate of which slope to follow. The scipy function odeint in Python uses an adaptive version of this internally, automatically shrinking $\Delta t$ in regions where the solution changes rapidly (e.g. near the epidemic peak) and enlarging it where the solution is flat.